3.b. Suppose there is a change in the demand of cold drinks from 700 Units to 1000 Units because of the fall in the price from Rs 15 to Rs 10. Calculate the price elasticity of cold drinks. 


1
Expert's answer
2021-05-11T14:25:15-0400

Given,

Initial demands =700 units

Final demands = 1000 units


Initial price = 15 Rs.

Final price = 10Rs.


Applying the arc elasticity,

Arc elasticity "=\dfrac{\frac{\frac{Q_2 -Q_1}{Q_2+Q_1}}{2}}{\dfrac{\frac{P_2-P_1}{P_2+P_1}}{2}}"

"=\frac{(Q_2-Q_1)(P_2+P_1)}{(Q_2+Q_1)(P_2-P_1)}"

Now, substituting the values,


"=-\frac{300\times 25}{1700\times 5}"


"=-\frac{15}{17}"

Hence the price elasticity demand = -15/17